# MEEG and AT-DGNN: Improving EEG Emotion Recognition with Music Introducing and Graph-based Learning

### v2

@@ Abstract
Neuropsychological research highlights the essential role of coordinated activities across brain regions during cognitive tasks. This study expands the existing EEG datasets by constructing the MEEG dataset, a multi-modal compilation of music-induced electroencephalogram (EEG) recordings, and further investigates the brain network topology during emotional responses to music. The MEEG dataset, capturing emotional responses to various musical stimuli across different valence and arousal levels, enables an in-depth analysis of brainwave patterns within musical contexts. We introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG emotion recognition. By integrating an attention mechanism with a dynamic graph neural network (DGNN), the AT-DGNN model captures complex local and global EEG dynamics, demonstrating superior performance with accuracy of 83.74% in arousal and 86.01% in valence, outperforming current state-of-the-art (SOTA) methods. Comparative analyses with traditional datasets like DEAP underscore our model’s effectiveness, highlighting the potential of music as a potent emotional stimulus. This study advances graph-based learning techniques in brain-computer interfaces (BCI), significantly enhancing the precision of EEG-based emotion recognition and deepening our understanding of cognitive functions in various brain regions. The source code and dataset are accessible at https://github.com/xmh1011/AT-DGNN.
@@ I Introduction
BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-machine interaction [3, 4]. This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands. Due to their objectivity, high temporal resolution, affordability, and rapid acquisition, EEG signals are indispensable for emotion recognition and neuroscience research [5, 6].
@@
In 2023, Ding et al. introduced the Local-Global-Graph network (LGGNet) to process EEG data in an image-like format, improving classification performance [16]. However, LGGNet’s single-layer graph neural network (GNN) struggles to capture the dynamic features of EEG signals. Additionally, popular EEG datasets like DEAP and SEED primarily use video stimuli, limiting the study of the relationship between music, emotions, and EEG signals. In contrast, EEG datasets and research using other types of stimuli are relatively scarce.
To address shortcomings in emotion recognition research, we create the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format with music-induced emotional states to enhance ecological validity. We improve the LGGNet architecture by integrating an attention mechanism and using a multi-layer dynamic graph neural network, significantly boosting classification accuracy on the MEEG dataset. This innovative approach allows for a more nuanced examination of brain function interconnections via graph-based learning.
@@
    We establish the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format, enhanced with various types of music to effectively induce emotional states. Experimental results show that the MEEG dataset outperforms the video-stimuli-based DEAP dataset in emotion induction, effectively enhancing model ACC and F1 scores.
@@
    We propose the AT-DGNN framework for exploring connections within and between different functional areas of the brain. By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion recognition.
@@
    We compare the performance of the proposed method with CNN, TCN, and GNN-based SOTA methods on the MEEG dataset. Extensive ablation experiments are conducted to better understand AT-DGNN.
@@ III-A EEG Data Preprocessing
The initial sampling rate of the MEEG dataset is 1000 Hz. To align with standard EEG datasets typically sampled at 128 Hz or 200 Hz, and to mitigate issues such as high artifact noise and overlapping interference signals, the data are downsampled to 200 Hz. This preprocessing phase involves the use of a band-pass filter with a range of 1 to 50 Hz to enhance the signal quality by minimizing interference. Subsequently, the EEG signals are segmented into five distinct frequency bands for feature extraction: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23].
@@ III-B1 Temporal Learner
A temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) is used to directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length. These kernels obviate the need for manually extracted features. The $i$th-kernel’s length ($S_{T}^{i}$), dictated by the sampling frequency ($f_{s}$) of EEG data and scaling coefficients ($\alpha^{i}$), is given by [24]:
@@ III-B2 Sliding Window Segmentation
Following the temporal learner, a convolution-based sliding window technique as [25] is applied to segment the EEG time series $X$ into multiple windows. This method integrates sliding window segmentation with convolutional operations, significantly reducing computational overhead by allowing convolutions to be executed once across all windows. This approach not only enhances data augmentation but also accelerates processing through parallel computation. The time series $X$ is segmented into windows $X_{w}\in\mathbb{R}^{B\times C\times W}$ using a window of length $W$ and stride $S$, where $w=1,\ldots,n$ represents the window index and $n$ is the total number of windows. Each window $X_{w}$ is then processed by subsequent attention and temporal convolution blocks. The number of windows $n$ is calculated as:
@@ III-B4 Temporal Convolution
To extract dynamic temporal features, we employ a temporal convolutional block on each window of the EEG data. The normalized EEG data $X_{i}$ serves as the input to this block. The temporal convolutional block processes $X_{i}$ and generates an output tensor $Z_{\textit{tcn}}^{w}\in\mathbb{R}^{B\times C\times T_{w}}$, where $T_{w}$ represents the window’s temporal length and $w$ indexes the window. Batch normalization is applied to the output to ensure training stability, followed by a $\textit{ReLU}(\cdot)$ activation function to introduce non-linearity. The output $Z_{\textit{tcn}}^{w}$ is computed as follows:
@@ III-C1 Graph Filtering Layer
For the extracted features, we adopt the method described by [16], which involves defining functional areas and performing local filtering. The electrodes are divided into three functional areas: the general region ($G_{g}$), the frontal region ($G_{f}$), and the hemispheric region ($G_{h}$), as shown in Fig. 2. Each channel is treated as a node, and the learned dynamic temporal representations are considered as node attributes.
Figure 2: Local global graph definitions [16]. (a) General definition $G_{g}$. (b) Frontal definition $G_{f}$. (c) Hemispheric definition $G_{h}$.
@@ III-C2 Stacked DGNN
21:do graph filtering and aggregation on each node by (12) - 15 to get $Z^{i}_{\textit{local}}$;
@@
We introduce a stacked DGNN architecture to discern complex structural relationships within graphs. Each layer adaptively recalculates the adjacency matrix based on input features, effectively capturing hierarchical feature variations.
@@ IV-A MEEG Dataset
In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments. However, participants’ responses to these stimuli may be influenced by their cultural backgrounds [26]. Considering that the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [27], we opted to induce emotions in participants using music in the MEEG dataset. We carefully selected 20 music clips, each lasting one minute, characterized by distinct arousal and valence levels, decidedly non-neutral. To minimize emotional carryover, a 15-second pause was introduced between clips [28]. The chosen music pieces included lesser-known works of Western classical music by composers such as Shostakovich and Tchaikovsky. The study involved 32 students from Shandong University, aged between 20 and 25 years, who had no formal music education and were unfamiliar with Western classical music. Each participant signed an informed consent form prior to the study. This experimental design aimed to minimize the influence of social and cultural backgrounds on emotional responses, ensuring that changes in emotions were solely due to musical stimuli, thereby enhancing the accuracy of the experimental results. EEG data were collected using a 32-channel BCI device NeuSen.W32 from Neuracle Tech, employing the same electrode channels as in the DEAP dataset [7]. The data were sampled at a frequency of 1000 Hz. The EEG data were annotated based on the arousal and valence of the music during various stages of the experiment.
@@ IV-B Experiment Settings
Additionally, a two-stage training strategy within the inner loop optimizes the utilization of training data. Initially, the best model identified from the $k$-fold cross-validation is saved as a preliminary candidate. This model is then refined using combined training data from the $k$ folds, fine-tuned at a reduced learning rate to prevent overfitting, and further trained for up to 20 epochs or until it achieves a training accuracy of 100%, ensuring precise calibration. Importantly, test data is excluded from the training process to maintain the integrity of the evaluation.
@@ IV-C Implement Details
The model was implemented using PyTorch [30] library, and the code can be found at https://github.com/xmh1011/AT-DGNN.
Cross-entropy loss was chosen as the objective function to guide the training process. The training was divided into two stages, with the first stage capped at 200 epochs and the second at 20 epochs. The batch size was set to 64. To reduce training time and prevent overfitting, early stopping was implemented. For the attention module, the window size was set to half the data frequency. The kernel sizes for the time learner were set to 100, 50, and 25. The hidden layer count for the DGNN was established at 32. Training was optimized using the Adam optimizer, starting with an initial learning rate of $31e-$ 3, which was reduced by a factor of 10 during the second stage. For more details, please refer to the open-source GitHub repository for AT-DGNN.
@@
LGGNet-Fro 81.85 12.57 81.58 12.73 82.34 12.84 81.66 13.91 82.11 81.62
LGGNet-Gen 82.15 12.72 82.15 12.57 82.53 13.20 82.47 13.22 82.34 82.31
LGGNet-Hem 81.92 12.08 81.58 12.43 82.37 12.71 81.58 14.13 82.15 81.58
@@ V-A Emotion Recognition
As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy. The AT-DGNN-Gen model not only achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, but also records the highest accuracy of 86.01% in the valence dimension, a substantial improvement of 3.48% over the LGGNet series. This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
The AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions. This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks. Furthermore, consistent with the findings of [16], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
Furthermore, these results demonstrate a marked improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy surpassing previous models by up to 7.16% and 5.78%, respectively. This emphasizes the advanced capability of the DGNN approach in graph-structured data analysis, providing more accurate and reliable emotion recognition.
All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%. On the DEAP dataset, AT-DGNN averages 60.56% accuracy across four dimensions. According to Huang et al. [32], other models, including LGGNet, also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not. All models perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
@@ V-B Ablation Study
As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of critical components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and the dynamic graph neural network (D). The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance. Modifying the number of DGNN layers to $1$, $2$, or $4$ led to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively. A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
The optimal configuration, employing $3$ layers of the DGNN, is found to be the most effective, as evidenced by achieving the highest scores in both ACC and F1 metrics. These findings substantiate the design choices of the AT-DGNN model and shed light on the interaction among its components, significantly enhancing computational neuroscience analysis.

### v3

@@ Abstract
We present the MEEG dataset, a multi-modal collection of music-induced electroencephalogram (EEG) recordings designed to capture emotional responses to various musical stimuli across different valence and arousal levels. This public dataset facilitates an in-depth examination of brainwave patterns within musical contexts, providing a robust foundation for studying brain network topology during emotional processing. Leveraging the MEEG dataset, we introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition. This model combines an attention mechanism with a dynamic graph neural network (DGNN) to capture intricate EEG dynamics. The AT-DGNN achieves state-of-the-art (SOTA) performance with an accuracy of 83.74% in arousal recognition and 86.01% in valence recognition, outperforming existing SOTA methods. Comparative analysis with traditional datasets, such as DEAP, further validates the model’s effectiveness and underscores the potency of music as an emotional stimulus. This study advances graph-based learning methodology in brain-computer interfaces (BCI), significantly improving the accuracy of EEG-based emotion recognition. The MEEG dataset and source code are publicly available at https://github.com/xmh1011/AT-DGNN.
@@ I Introduction
BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-computer interaction [3]. This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands. Due to their objectivity, high temporal resolution, affordability, and rapid acquisition, EEG signals are indispensable for emotion recognition and neuroscience research [4, 5].
@@
In 2023, Ding et al. introduced the Local-Global-Graph Network (LGGNet) for processing EEG data in an image-like format, which improved classification performance [15]. Alternatively, LGGNet’s single-layer graph neural network faces challenges in capturing the dynamic features of EEG signals. Additionally, widely used EEG datasets such as DEAP and SEED primarily involve visual stimuli, which limits exploration of the relationship between music, emotions, and EEG signals. Conversely, EEG datasets and research using musical stimuli are relatively scarce. Moreover, existing SOTA methods achieve lower accuracy on the DEAP dataset, thus limiting comprehensive performance evaluation of models.
To address these limitations in emotion recognition research, the MEEG (Music EEG) dataset is developed accordingly. This multi-modal dataset, similar to DEAP, utilizes music-induced emotional states to improve the accuracy of emotion recognition. By integrating a sliding window technique, an attention mechanism, and a multi-layer dynamic graph neural network into the LGGNet architecture, significant improvements in classification accuracy are achieved.
@@
    The MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format, is enhanced with diverse music to induce emotional states effectively. It outperforms the DEAP dataset in emotion induction, improving model accuracy (ACC) and F1 scores.
@@
    The AT-DGNN framework explores connections within and between brain functional areas. By integrating a sliding window technique, an attention mechanism, and a stacked DGNN, this novel architecture for emotion recognition enhances the learning of dynamic features in EEG recordings.
@@
    Performance of the AT-DGNN is compared with other CNN, TCN, and GNN-based SOTA methods on MEEG dataset. Ablation experiments provide insights into the AT-DGNN framework.
@@ III-A EEG Data Preprocessing
The initial sampling rate of the MEEG dataset is 1000 Hz. To align with standard datasets typically sampled at 128 Hz or 200 Hz, and to mitigate issues such as high artifact noise and overlapping interference signals, the data are downsampled to 200 Hz. This preprocessing phase involves the use of a band-pass filter with a range of 1 to 50 Hz to enhance the signal quality by minimizing interference. Subsequently, the EEG signals are segmented into five distinct frequency bands for feature extraction: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz).
@@ III-B1 Temporal Learner
A temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) is used to directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length. These kernels obviate the need for manually extracted features. The $i$-th kernel’s length ($S_{T}^{i}$), dictated by the sampling frequency ($f_{s}$) of EEG data and scaling coefficients ($\alpha^{i}$), can be represented by [22]:
@@ III-B2 Sliding Window Segmentation
Following the temporal learner, a convolution-based sliding window technique as described in [23] is applied to segment the EEG time series $X$ into multiple windows. This method integrates sliding window segmentation with convolutional operations, significantly reducing computational overhead by allowing convolutions to be executed once across all windows. This approach enhances data augmentation while accelerating processing through parallel computation. The time series $X$ is segmented into windows $X_{w}\in\mathbb{R}^{B\times C\times W}$ using a window of length $W$ and stride $S$, where $w=1,\ldots,n$ represents the window index and $n$ is the total number of windows. Each window $X_{w}$ is then processed by subsequent attention and temporal convolution blocks. The number of windows $n$ is calculated as:
@@ III-B4 Temporal Convolution
To extract dynamic temporal features, a temporal convolutional block is applied to each window of the EEG data. The normalized EEG data $X_{i}$ serves as the input to this block. The temporal convolutional block processes $X_{i}$ to produce an output tensor $Z_{\textit{tcn}}^{w}\in\mathbb{R}^{B\times C\times T_{w}}$, where $T_{w}$ is the temporal length of the window and $w$ is the window index. Batch normalization is applied to ensure training stability, followed by a $\textit{ReLU}(\cdot)$ activation function to introduce non-linearity. The output $Z_{\textit{tcn}}^{w}$ is computed as follows:
@@ III-C1 Graph Filtering Layer
The method described by [15] is adopted for the extracted features, involving the definition of functional areas and local filtering. Electrodes are divided into three functional areas: the general region ($G_{g}$), the frontal region ($G_{f}$), and the hemispheric region ($G_{h}$), as shown in Fig. 2. Each channel is treated as a node, with the learned dynamic temporal representations considered as node attributes.
Figure 2: Local-global graph definitions [15]. (a) General definition $G_{g}$. (b) Frontal definition $G_{f}$. (c) Hemispheric definition $G_{h}$.
@@ III-C2 Stacked DGNN
21:perform graph filtering and aggregation on each node by (12) - (15) to get $Z^{i}_{\textit{local}}$;
@@
A stacked Directed Graph Neural Network (DGNN) captures complex graph relationships by recalculating the adjacency matrix at each layer based on input features, effectively capturing hierarchical feature variations.
@@ IV-A MEEG Dataset
In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments. However, participants’ responses to these stimuli may be influenced by their cultural backgrounds [24]. Considering that the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [25], we opted to induce emotions in participants using music in the MEEG dataset. Under the guidance of music professors, we selected 20 lesser-known Western classical music clips by composers such as Shostakovich and Tchaikovsky. Each clip lasts one minute and is characterized by distinct levels of valence and arousal. To minimize emotional carryover, a 15-second pause was introduced between clips [26]. The study involved 32 students from Shandong University, aged between 20 and 25 years, who had no formal music education and were unfamiliar with Western classical music. Each participant signed an informed consent form prior to the study. This experimental design aimed to minimize the influence of social and cultural backgrounds on emotional responses, ensuring that changes in emotions were solely due to musical stimuli, thereby enhancing the accuracy of the experimental results. EEG data were collected using a 32-channel BCI device NeuSen.W32 from Neuracle Tech, employing the same electrode channels as in the DEAP dataset [6]. The data were sampled at a frequency of 1000 Hz. The EEG data were annotated based on the arousal and valence of the music during various stages of the experiment.
In contrast to the DEAP dataset, the MEEG dataset reduces experimental bias by rigorously selecting participants and musical stimuli, minimizing subjective factors and improving the effectiveness of emotional elicitation.
@@ IV-B Experiment Settings
Additionally, a two-stage training strategy within the inner loop optimizes the utilization of training data. Initially, the best model identified from the $k$-fold cross-validation is saved as a preliminary candidate. This model is subsequently refined with the aggregated data from all $k$ folds, fine-tuned at a lower learning rate to avoid overfitting, and further trained for a maximum of 20 epochs or until it reaches 100% training accuracy, ensuring precise calibration. Importantly, test data is excluded from the training to preserve evaluation integrity.
@@ IV-C Implement Details
The model was implemented using PyTorch [28] library.
Cross-entropy loss was chosen as the objective function to guide the training process. The training was divided into two stages, with the first stage capped at 200 epochs and the second at 20 epochs. To reduce training time and prevent overfitting, early stopping was implemented. For the attention module, the window size was set to half the $f_{s}$. The kernel sizes for the time learner were set to 100, 50, and 25. Training was optimized using the Adam optimizer, starting with an initial learning rate of $31e-$ 3, which was reduced by a factor of 10 during the second stage. For more information on the data processing, model implementation details, and accessing the dataset, please refer to our GitHub repository https://github.com/xmh1011/AT-DGNN.
@@
LGGNet-Fro 81.85 12.57 81.58 12.73 84.28 11.83 83.87 12.66 83.07 82.73
LGGNet-Gen 82.15 12.72 82.15 12.57 84.53 11.44 84.38 11.56 83.34 83.27
LGGNet-Hem 81.92 12.08 81.58 12.43 84.93 11.21 84.46 12.27 83.42 83.02
@@ V-A Emotion Recognition
As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy. The AT-DGNN-Gen model achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, and also records the highest accuracy of 86.01% in the valence dimension, marking a substantial improvement of 1.08% over the LGGNet series. This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
The AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions. This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks. Furthermore, consistent with the findings of LGGNet [15], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
Furthermore, the results show a significant improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy exceeding previous models by up to 7.16% and 5.78%, respectively. This highlights the enhanced capability of the DGNN approach in analyzing graph-structured data, offering more accurate and reliable emotion recognition. Additionally, the AT-DGNN exhibits a lower standard deviation compared to other models, indicating more balanced performance across various samples and superior generalization ability.
All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%. On the DEAP dataset, AT-DGNN achieves an average accuracy of 60.56% across four dimensions, while LGGNet reaches 60.96%. According to Huang et al. [29], other SOTA methods also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not. All methods perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
@@ V-B Ablation Study
As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of key components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and dynamic graph neural network (D). The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance. Modifying the number of DGNN layers to $1$, $2$, or $4$ lead to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively. A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
The optimal configuration, utilizing $3$ layers of the DGNN, demonstrated superior effectiveness by achieving the highest scores in both ACC and F1 metrics, substantiating the design choices and highlighting the critical interdependence of its three main components: sliding window segmentation, multi-head attention mechanisms, and dynamic graph neural networks. These findings reveal how the synergistic operation of these elements is essential for capturing the dynamic temporal features of EEG data, significantly enhancing the model’s capability to process EEG data effectively and advancing computational neuroscience analysis.